calculate the number of moles in 200grams (1)o2 (2)co2 (3)N2(4)NH3(5)H2
1
Expert's answer
2014-11-03T12:02:29-0500
1) The molar mass of O2is 32g/mole. So, the number of moles in 200g of O2 = 200g/32g/mole = 6.25 moles.
2) The molar mass of CO2 is 44g/mole. So, the number of moles in 200g of CO2 is 200g/44g/mole = 4.54 moles.
3) The molar mass of N2is 28g/mole. So, the number of moles in 200g of N2 is 200g/28g/mole = 7.14 moles.
4) The molar mass ofNH3 is 17g/mole. So, the number of moles in 200g of NH3 is 200g/17g/mole = 11.76 moles.
5) The molar mass of H2 is 2g/mole. So, the number of moles in 200g of H2 is 200g/2g/mole = 100 moles.